What is the conditional probability P(B|A) of dealing a full house consisting of 3 aces and 2 kings if we are told by the dealer before he gives us our five cards that our hand has at least 1 ace in it already.
Answer #1

total cards=52,

total ace cards=4, remaining cards = 48

P(at least one ace) = 1 - P(no ace) = 1 - 48C5/52C5 = 0.3412

P(3 aces and 2 kings | atleast one ace) = (4C3*4C2/52C5) /0.3412 = 0.000027068